OpenStudy (anonymous):

What's the limit as x approaches infinity, (5x+8)/(10x^2+5x+2)?

OpenStudy (jamesj):

Divide top and bottom by x ... then what do you think the answer is intuitively?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[\lim{_x{\rightarrow}}infinity\frac{5+\frac{8}{x}}{10x + 5 +\frac{2}{x}}\]

OpenStudy (anonymous):

ok, i got that and then i was able to deduce that it simplified to \[5\div(10x-5)\]

OpenStudy (anonymous):

well i meant + 5

OpenStudy (jamesj):

Then by the same logic this ratio also goes to zero as x goes to infinity

OpenStudy (anonymous):

oh, ok...thank ya for clarifying that